# High-Strength Amorphous Silicon Carbide for Nanomechanics

### v1

@@ VI Conclusion and outlook
In summary, our study has uncovered an amorphous silicon carbide thin film with a ultimate tensile strength above 10 GPa, the highest value ever measured for a nanostructured amorphous material and approaching the experimental values shown by graphene nano-ribbons [19]. Their robustness to chemicals allow us to fabricate nanostructures with very high fidelity even when their geometries make them delicate high-aspect-ratio structures. This ability to produce structures with high fidelity also allow us to measure the film’s mechanical properties with high precision. We deposit amorphous silicon carbide in varying deposition conditions and substrates to understand new approaches towards increasing ultimate yield strength. Then using the a-SiC with the highest UTS, we designed and fabricated a variety of well-understood nanostructures such as cantilever, membranes and doubly-clamped strings to measure the thin films mechanical properties such as density, Young’s modulus, Poisson ratio, and mechanical loss tangents. For the latter we employ nanostrings patterned with phononic bandstructures which conventionally have some of the lowest mechanical dissipations in literature, and this allows us to measure very low mechanical dissipation. The a-SiC nanostrings support soft-clamped mechanical modes with quality factors exceeding at room temperature; a new regime for SiC devices and on par with the state-of-the-art SiN resonators. This corresponds to a high force sensitivity of 7.7 aN/Hz. We demonstrate a robust characterization process based on the simple fabrication and optical techniques which does not rely on complex tension loading setups.
@@ Acknowledgements.
We wish to acknowledge Peter G. Steeneken, Gerard Verbiest, and Martin Lee for their helpful suggestions on the manuscript and their support of our project. We want to thank Satadal Dutta, Ali Sarafraz, Matthijs de Jong for the helpful discussions. M.X. and R.N. also thank the staffs of both the Kavli Nanolab Delft and the Else Kooi Lab, in particular from C. de Boer, for supporting our fabrication efforts. This publication is part of the project, Probing the physics of exotic superconductors with microchip Casimir experiments (740.018.020) of the research programme NWO Start-up which is partly financed by the Dutch Research Council (NWO).This work has received funding from the EMPIR programme co-financed by the Participating States and from the European Union’s Horizon 2020 research and innovation programme (No. 17FUN05 PhotoQuant). R.N. would like to acknowledge support from the Limitless Space Institute’s I Grant.References
- Krause et al. [2012] A. G. Krause, M. Winger, T. D. Blasius, Q. Lin, and O. Painter, A high-resolution microchip optomechanical accelerometer, Nature Photonics 6, 768 (2012).
@@
- Pratt et al. [2023] J. R. Pratt, A. R. Agrawal, C. A. Condos, C. M. Pluchar, S. Schlamminger, and D. J. Wilson, Nanoscale torsional dissipation dilution for quantum experiments and precision measurement, Physical Review X 13, 011018 (2023).
- Manzaneque et al. [2022] T. Manzaneque, M. K. Ghatkesar, F. Alijani, M. Xu, R. A. Norte, and P. G. Steeneken, Resolution limits of resonant sensors with duffing non-linearity, arXiv preprint arXiv:2205.11903 (2022).
@@
- Garcia-Sanchez et al. [2012] D. Garcia-Sanchez, K. Y. Fong, H. Bhaskaran, S. Lamoreaux, and H. X. Tang, Casimir force and in situ surface potential measurements on nanomembranes, Physical review letters 109, 027202 (2012).
- Pate et al. [2020] J. M. Pate, M. Goryachev, R. Y. Chiao, J. E. Sharping, and M. E. Tobar, Casimir spring and dilution in macroscopic cavity optomechanics, Nature Physics 16, 1117 (2020).
@@
- Hälg et al. [2021] D. Hälg, T. Gisler, Y. Tsaturyan, L. Catalini, U. Grob, M.-D. Krass, M. Héritier, H. Mattiat, A.-K. Thamm, R. Schirhagl, et al., Membrane-based scanning force microscopy, Physical Review Applied 15, L021001 (2021).
- Reinhardt et al. [2016] C. Reinhardt, T. Müller, A. Bourassa, and J. C. Sankey, Ultralow-noise sin trampoline resonators for sensing and optomechanics, Physical Review X 6, 021001 (2016).
@@
- Verbridge et al. [2006] S. S. Verbridge, J. M. Parpia, R. B. Reichenbach, L. M. Bellan, and H. G. Craighead, High quality factor resonance at room temperature with nanostrings under high tensile stress, Journal of Applied Physics 99, 124304 (2006).
- Norte et al. [2016] R. A. Norte, J. P. Moura, and S. Gröblacher, Mechanical resonators for quantum optomechanics experiments at room temperature, Physical review letters 116, 147202 (2016).
@@
- Ghadimi et al. [2018] A. H. Ghadimi, S. A. Fedorov, N. J. Engelsen, M. J. Bereyhi, R. Schilling, D. J. Wilson, and T. J. Kippenberg, Elastic strain engineering for ultralow mechanical dissipation, Science 360, 764 (2018).
- Fedorov et al. [2020] S. A. Fedorov, A. Beccari, N. J. Engelsen, and T. J. Kippenberg, Fractal-like mechanical resonators with a soft-clamped fundamental mode, Physical Review Letters 124, 025502 (2020).
- Shin et al. [2022] D. Shin, A. Cupertino, M. H. de Jong, P. G. Steeneken, M. A. Bessa, and R. A. Norte, Spiderweb nanomechanical resonators via bayesian optimization: inspired by nature and guided by machine learning, Advanced Materials 34, 2106248 (2022).
@@
- Bereyhi et al. [2022] M. J. Bereyhi, A. Arabmoheghi, A. Beccari, S. A. Fedorov, G. Huang, T. J. Kippenberg, and N. J. Engelsen, Perimeter modes of nanomechanical resonators exhibit quality factors exceeding at room temperature, Physical Review X 12, 021036 (2022).
@@
- Beccari et al. [2022] A. Beccari, D. A. Visani, S. A. Fedorov, M. J. Bereyhi, V. Boureau, N. J. Engelsen, and T. J. Kippenberg, Strained crystalline nanomechanical resonators with quality factors above 10 billion, Nature Physics 18, 436 (2022).
@@
- Romero et al. [2020] E. Romero, V. M. Valenzuela, A. R. Kermany, L. Sementilli, F. Iacopi, and W. P. Bowen, Engineering the dissipation of crystalline micromechanical resonators, Physical Review Applied 13, 044007 (2020).
- Klaß [2022] Y. S. Klaß, High Q nanomechanical resonators fabricated from crystalline silicon carbide, Ph.D. thesis, Technische Universität München (2022).
@@
- Manjeshwar et al. [2022] S. K. Manjeshwar, A. Ciers, F. Hellman, J. Bläsing, A. Strittmater, and W. Wieczorek, Micromechanical high-q trampoline resonators from strained crystalline ingap for integrated free-space optomechanics, arXiv preprint arXiv:2211.12469 (2022).
- Wu et al. [2021] Q. Wu, X. Yan, Y. Jia, and X. Yao, Defective carbon-based materials: controllable synthesis and electrochemical applications, EnergyChem 3, 100059 (2021).
- Goldsche et al. [2018] M. Goldsche, J. Sonntag, T. Khodkov, G. J. Verbiest, S. Reichardt, C. Neumann, T. Ouaj, N. von den Driesch, D. Buca, and C. Stampfer, Tailoring mechanically tunable strain fields in graphene, Nano letters 18, 1707 (2018).
- Cao et al. [2020] K. Cao, S. Feng, Y. Han, L. Gao, T. Hue Ly, Z. Xu, and Y. Lu, Elastic straining of free-standing monolayer graphene, Nature communications 11, 284 (2020).
- Gu et al. [2013] X. W. Gu, Z. Wu, Y.-W. Zhang, D. J. Srolovitz, and J. R. Greer, Microstructure versus flaw: mechanisms of failure and strength in nanostructures, Nano letters 13, 5703 (2013).
- Zhao et al. [2022] X. Zhao, B. Mao, M. Liu, J. Cao, S. J. Haigh, D. G. Papageorgiou, Z. Li, and R. J. Young, Controlling and monitoring crack propagation in monolayer graphene single crystals, Advanced Functional Materials 32, 2202373 (2022).
@@
- Bereyhi et al. [2019] M. J. Bereyhi, A. Beccari, S. A. Fedorov, A. H. Ghadimi, R. Schilling, D. J. Wilson, N. J. Engelsen, and T. J. Kippenberg, Clamp-tapering increases the quality factor of stressed nanobeams, Nano letters 19, 2329 (2019).
- Kwon et al. [2015] G. Kwon, H.-H. Jo, S. Lim, C. Shin, H.-H. Jin, J. Kwon, and G.-M. Sun, Room-temperature yield and fracture strength of single-crystalline 6h silicon carbide, Journal of materials science 50, 8104 (2015).
- Rasool et al. [2013] H. I. Rasool, C. Ophus, W. S. Klug, A. Zettl, and J. K. Gimzewski, Measurement of the intrinsic strength of crystalline and polycrystalline graphene, Nature communications 4, 2811 (2013).
- Zhang et al. [2015] T. Zhang, X. Li, and H. Gao, Fracture of graphene: a review, International Journal of Fracture 196, 1 (2015).
- Lee et al. [2008] C. Lee, X. Wei, J. W. Kysar, and J. Hone, Measurement of the elastic properties and intrinsic strength of monolayer graphene, science 321, 385 (2008).
- Wang et al. [2012] M. Wang, C. Yan, L. Ma, N. Hu, and M. Chen, Effect of defects on fracture strength of graphene sheets, Computational Materials Science 54, 236 (2012).
- Xu et al. [2018] J. Xu, G. Yuan, Q. Zhu, J. Wang, S. Tang, and L. Gao, Enhancing the strength of graphene by a denser grain boundary, ACS nano 12, 4529 (2018).
- Qu et al. [2012] R. Qu, M. Calin, J. Eckert, and Z. Zhang, Metallic glasses: Notch-insensitive materials, Scripta Materialia 66, 733 (2012).
- Jiang et al. [2021] H. Jiang, T. Shang, H. Xian, B. Sun, Q. Zhang, Q. Yu, H. Bai, L. Gu, and W. Wang, Structures and functional properties of amorphous alloys, Small Structures 2, 2000057 (2021).
- Greer [1995] A. L. Greer, Metallic glasses, Science 267, 1947 (1995).
- Telford [2004] M. Telford, The case for bulk metallic glass, Materials today 7, 36 (2004).
- Wong et al. [1997] E. W. Wong, P. E. Sheehan, and C. M. Lieber, Nanobeam mechanics: elasticity, strength, and toughness of nanorods and nanotubes, science 277, 1971 (1997).
- Cui et al. [2019] J. Cui, Z. Zhang, H. Jiang, D. Liu, L. Zou, X. Guo, Y. Lu, I. P. Parkin, and D. Guo, Ultrahigh recovery of fracture strength on mismatched fractured amorphous surfaces of silicon carbide, ACS nano 13, 7483 (2019).
@@
- Wijesundara and Azevedo [2011] M. Wijesundara and R. Azevedo, Silicon carbide microsystems for harsh environments, Vol. 22 (Springer Science & Business Media, 2011).
- Gerhardt [2011] R. Gerhardt, Properties and applications of silicon carbide (BoD–Books on Demand, 2011).
@@
- Kimoto and Cooper [2014] T. Kimoto and J. A. Cooper, Fundamentals of Silicon Carbide Technology: Growth, Characterization, Devices and Applications (Wiley-IEEE Press, 2014).
- Yi et al. [2020] A. Yi, Y. Zheng, H. Huang, J. Lin, Y. Yan, T. You, K. Huang, S. Zhang, C. Shen, M. Zhou, et al., Wafer-scale 4h-silicon carbide-on-insulator (4h–sicoi) platform for nonlinear integrated optical devices, Optical Materials 107, 109990 (2020).
- Martini and Politi [2017] F. Martini and A. Politi, Linear integrated optics in 3c silicon carbide, Optics Express 25, 10735 (2017).
- Powell et al. [2022] K. Powell, L. Li, A. Shams-Ansari, J. Wang, D. Meng, N. Sinclair, J. Deng, M. Lončar, and X. Yi, Integrated silicon carbide electro-optic modulator, Nature Communications 13, 1851 (2022).
- Saddow and Agarwal [2004] S. E. Saddow and A. K. Agarwal, Advances in silicon carbide processing and applications (Artech House, 2004).
- Morana et al. [2013] B. Morana, G. Pandraud, J. Creemer, and P. Sarro, Characterization of lpcvd amorphous silicon carbide (a-sic) as material for electron transparent windows, Materials Chemistry and Physics 139, 654 (2013).
@@
- Iliescu and Poenar [2012] C. Iliescu and D. P. Poenar, Pecvd amorphous silicon carbide (-sic) layers for mems applications, in Physics and Technology of Silicon Carbide Devices (IntechOpen, 2012).
- Fraga et al. [2012] M. A. Fraga, R. S. Pessoa, M. Massi, and H. S. Maciel, Applications of sic-based thin films in electronic and mems devices, Physics and technology of silicon carbide devices 1, 313 (2012).
- Blank et al. [1999] V. Blank, M. Popov, G. Pivovarov, N. Lvova, and S. Terentev, Mechanical properties of different types of diamond, Diamond and related materials 8, 1531 (1999).
- Dang et al. [2021] C. Dang, J.-P. Chou, B. Dai, C.-T. Chou, Y. Yang, R. Fan, W. Lin, F. Meng, A. Hu, J. Zhu, et al., Achieving large uniform tensile elasticity in microfabricated diamond, Science 371, 76 (2021).
- Papageorgiou et al. [2017] D. G. Papageorgiou, I. A. Kinloch, and R. J. Young, Mechanical properties of graphene and graphene-based nanocomposites, Progress in materials science 90, 75 (2017).
- Blum et al. [1999] T. Blum, B. Dresler, M. Hoffmann, et al., Wear-resistant amorphous sic coatings produced by plasma-enhanced cvd, Surface and Coatings Technology 116, 1024 (1999).
- Jiang et al. [1999] L. Jiang, X. Chen, X. Wang, L. Xu, F. Stubhan, and K.-H. Merkel, a-sicx: H films deposited by plasma-enhanced chemical vapor deposition at low temperature used for moisture and corrosion resistant applications, Thin Solid Films 352, 97 (1999).
- Flannery et al. [1998] A. F. Flannery, N. J. Mourlas, C. W. Storment, S. Tsai, S. H. Tan, J. Heck, D. Monk, T. Kim, B. Gogoi, and G. T. Kovacs, Pecvd silicon carbide as a chemically resistant material for micromachined transducers, Sensors and Actuators A: Physical 70, 48 (1998).
- Buijtendorp et al. [2022] B. Buijtendorp, S. Vollebregt, K. Karatsu, D. Thoen, V. Murugesan, K. Kouwenhoven, S. Hähnle, J. Baselmans, and A. Endo, Hydrogenated amorphous silicon carbide: A low-loss deposited dielectric for microwave to submillimeter-wave superconducting circuits, Physical Review Applied 18, 064003 (2022).
@@
- Köhler et al. [2021] M. Köhler, M. Pomaska, P. Procel, R. Santbergen, A. Zamchiy, B. Macco, A. Lambertz, W. Duan, P. Cao, B. Klingebiel, et al., A silicon carbide-based highly transparent passivating contact for crystalline silicon solar cells approaching efficiencies of 24%, Nature Energy 6, 529 (2021).
- Tawada et al. [1982] Y. Tawada, M. Kondo, H. Okamoto, and Y. Hamakawa, Hydrogenated amorphous silicon carbide as a window material for high efficiency a-si solar cells, Solar Energy Materials 6, 299 (1982).
- van Cleef et al. [1998] M. W. van Cleef, F. A. Rubinelli, R. Rizzoli, R. Pinghini, R. E. Schropp, and W. F. van der Weg, Amorphous silicon carbide/crystalline silicon heterojunction solar cells: a comprehensive study of the photocarrier collection, Japanese journal of applied physics 37, 3926 (1998).
- Sarro [2000] P. M. Sarro, Silicon carbide as a new mems technology, Sensors and Actuators A: Physical 82, 210 (2000).
- Fraga and Pessoa [2020] M. Fraga and R. Pessoa, Progresses in synthesis and application of sic films: From cvd to ald and from mems to nems, Micromachines 11, 799 (2020).
- Barnes et al. [2012] A. C. Barnes, C. A. Zorman, and P. X. Feng, Amorphous silicon carbide (-sic) thin square membranes for resonant micromechanical devices, in Materials Science Forum, Vol. 717 (Trans Tech Publ, 2012) pp. 533–536.
- Xing et al. [2019] P. Xing, D. Ma, K. J. Ooi, J. W. Choi, A. M. Agarwal, and D. Tan, Cmos-compatible pecvd silicon carbide platform for linear and nonlinear optics, ACS Photonics 6, 1162 (2019).
- Lee et al. [2015] J. Y. Lee, X. Lu, and Q. Lin, High-q silicon carbide photonic-crystal cavities, Applied Physics Letters 106, 041106 (2015).
- Xing et al. [2020] P. Xing, D. Ma, L. C. Kimerling, A. M. Agarwal, and D. T. Tan, High efficiency four wave mixing and optical bistability in amorphous silicon carbide ring resonators, APL Photonics 5, 076110 (2020).
- Auditore et al. [2002] A. Auditore, C. Satriano, U. Coscia, G. Ambrosone, V. Parisi, and G. Marletta, Human serum albumin adsorption onto a-sic: H and ac: H thin films deposited by plasma enhanced chemical vapor deposition, Biomolecular Engineering 19, 85 (2002).
- Saddow [2016] S. Saddow, Silicon Carbide Biotechnology, Second Edition: A Biocompatible Semiconductor for Advanced Biomedical Devices and Applications (Elsevier, 2016).
- Bauser et al. [2022] H. C. Bauser, M. D. Foley, M. E. Phelan, W. Weigand, D. R. Needell, Z. C. Holman, and H. A. Atwater, Amorphous silicon carbide high contrast gratings as highly efficient spectrally selective visible reflectors, Optics Express 30, 26787 (2022).
- Deku et al. [2018] F. Deku, C. L. Frewin, A. Stiller, Y. Cohen, S. Aqeel, A. Joshi-Imre, B. Black, T. J. Gardner, J. J. Pancrazio, and S. F. Cogan, Amorphous silicon carbide platform for next generation penetrating neural interface designs, Micromachines 9, 480 (2018).
- Teufel et al. [2009] J. D. Teufel, T. Donner, M. Castellanos-Beltran, J. W. Harlow, and K. W. Lehnert, Nanomechanical motion measured with an imprecision below that at the standard quantum limit, Nature nanotechnology 4, 820 (2009).
- Bagci et al. [2014] T. Bagci, A. Simonsen, S. Schmid, L. G. Villanueva, E. Zeuthen, J. Appel, J. M. Taylor, A. Sørensen, K. Usami, A. Schliesser, et al., Optical detection of radio waves through a nanomechanical transducer, Nature 507, 81 (2014).
- Castelletto and Boretti [2020] S. Castelletto and A. Boretti, Silicon carbide color centers for quantum applications, Journal of Physics: Photonics 2, 022001 (2020).
- Castelletto et al. [2022] S. Castelletto, A. Peruzzo, C. Bonato, B. C. Johnson, M. Radulaski, H. Ou, F. Kaiser, and J. Wrachtrup, Silicon carbide photonics bridging quantum technology, ACS Photonics 9, 1434 (2022).
- Sementilli et al. [2022] L. Sementilli, E. Romero, and W. P. Bowen, Nanomechanical dissipation and strain engineering, Advanced Functional Materials 32, 2105247 (2022).
@@
- Nguyen et al. [2017] T.-K. Nguyen, H.-P. Phan, H. Kamble, R. Vadivelu, T. Dinh, A. Iacopi, G. Walker, L. Hold, N.-T. Nguyen, and D. V. Dao, Superior robust ultrathin single-crystalline silicon carbide membrane as a versatile platform for biological applications, ACS applied materials & interfaces 9, 41641 (2017).
@@
- Atwater et al. [2018] H. A. Atwater, A. R. Davoyan, O. Ilic, D. Jariwala, M. C. Sherrott, C. M. Went, W. S. Whitney, and J. Wong, Materials challenges for the starshot lightsail, Nature materials 17, 861 (2018).
- Morana [2015] B. Morana, Silicon carbide thin films for MEMS nanoreactors for in-situ transmission electron microscopy, Ph.D. thesis, TU Delft (2015).
- Chu and Zhang [2009] J. Chu and D. Zhang, Mechanical characterization of thermal sio2 micro-beams through tensile testing, Journal of Micromechanics and Microengineering 19, 095020 (2009).
- Imran et al. [2018] M. Imran, M. Mahendran, and P. Keerthan, Mechanical properties of cold-formed steel tubular sections at elevated temperatures, Journal of Constructional Steel Research 143, 131 (2018).
- Zhou et al. [2004] Y. Zhou, Y. Wang, and P. Mallick, An experimental study on the tensile behavior of kevlar fiber reinforced aluminum laminates at high strain rates, Materials Science and Engineering: A 381, 355 (2004).
- Chen et al. [2000] K.-S. Chen, A. Ayon, and S. M. Spearing, Controlling and testing the fracture strength of silicon on the mesoscale, Journal of the American Ceramic Society 83, 1476 (2000).
- Shuman et al. [2007] D. J. Shuman, A. L. Costa, and M. S. Andrade, Calculating the elastic modulus from nanoindentation and microindentation reload curves, Materials characterization 58, 380 (2007).
- Kim et al. [2003] J.-H. Kim, S.-C. Yeon, Y.-K. Jeon, J.-G. Kim, and Y.-H. Kim, Nano-indentation method for the measurement of the poisson’s ratio of mems thin films, Sensors and Actuators A: Physical 108, 20 (2003).
- Klaß et al. [2022] Y. S. Klaß, J. Doster, M. Bückle, R. Braive, and E. M. Weig, Determining young’s modulus via the eigenmode spectrum of a nanomechanical string resonator, Applied Physics Letters 121, 083501 (2022).

### v2

@@ VI Conclusion and outlook
In summary, our study has uncovered an amorphous silicon carbide thin film with a ultimate tensile strength above 10 GPa, the highest value ever measured for a nanostructured amorphous material and approaching the experimental values shown by graphene nano-ribbons [23]. Their robustness to chemicals allow us to fabricate nanostructures with very high fidelity even when their geometries make them delicate high-aspect-ratio structures. This ability to reliably produce structures also allow us to measure the film’s mechanical properties with high precision. We deposit amorphous silicon carbide in varying deposition conditions and substrates to understand new approaches towards increasing ultimate yield strength. Then using the a-SiC with the highest UTS, we designed and fabricated a variety of well-understood nanostructures such as cantilevers, membranes and doubly-clamped strings to measure the thin films mechanical properties such as density, Young’s modulus, Poisson ratio, and mechanical loss tangent. For the latter we employ nanostrings patterned with phononic bandstructures which conventionally show some of the lowest mechanical dissipations in literature, and this allows us to measure very low mechanical dissipation. The a-SiC nanostrings support soft-clamped mechanical modes with quality factors exceeding at room temperature; a new regime for SiC devices and on par with the state-of-the-art SiN resonators. This corresponds to a high force sensitivity of 7.7 aN/Hz1/2. We demonstrate a robust characterization process based on the simple fabrication and optical techniques which does not rely on complex tension loading setups.
From molecular analysis of chemical composition (shown in the Supporting Information (H)), it appears that the elevated tensile strength in our a-SiC films is likely attributed to a higher prevalence of robust C-C bonds compared to weaker Si-Si bonds. This insight not only sheds light on the intricate interplay of bond structures and deposition parameters in determining mechanical resilience but also lays an understanding for further exploration in harnessing the unique properties of amorphous materials.
@@ Acknowledgements.
We wish to acknowledge Peter G. Steeneken, Gerard Verbiest, and Martin Lee for their helpful suggestions on the manuscript and their support of our project. We want to thank Satadal Dutta, Ali Sarafraz, Matthijs de Jong, Hanqing Liu for the helpful discussions. M.X. and R.N. also thank the staffs of both the Kavli Nanolab Delft and the Else Kooi Lab, in particular from Charles de Boer, for supporting our fabrication efforts, and from Hozanna Miro, for helping to perform the XRD measurement. The authours appreciate the help from P.R. Anusuyadevi and P. Gonugunta on the XPS measurement. This publication is part of the project, Probing the physics of exotic superconductors with microchip Casimir experiments (740.018.020) of the research programme NWO Start-up which is partly financed by the Dutch Research Council (NWO).This work has received funding from the EMPIR programme co-financed by the Participating States and from the European Union’s Horizon 2020 research and innovation programme (No. 17FUN05 PhotoQuant). R.N. would like to acknowledge support from the Limitless Space Institute’s I2 Grant. Funded/Co-funded by the European Union (ERC, EARS, 101042855). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.References
- [1] A. G. Krause, M. Winger, T. D. Blasius, Q. Lin, O. Painter, Nature Photonics 2012, 6, 11 768.
@@
- [2] J. R. Pratt, A. R. Agrawal, C. A. Condos, C. M. Pluchar, S. Schlamminger, D. J. Wilson, Physical Review X 2023, 13, 1 011018.
- [3] T. Manzaneque, M. K. Ghatkesar, F. Alijani, M. Xu, R. A. Norte, P. G. Steeneken, arXiv preprint arXiv:2205.11903 2022.
@@
- [4] D. Garcia-Sanchez, K. Y. Fong, H. Bhaskaran, S. Lamoreaux, H. X. Tang, Physical review letters 2012, 109, 2 027202.
- [5] J. M. Pate, M. Goryachev, R. Y. Chiao, J. E. Sharping, M. E. Tobar, Nature Physics 2020, 16, 11 1117.
@@
- [6] D. Hälg, T. Gisler, Y. Tsaturyan, L. Catalini, U. Grob, M.-D. Krass, M. Héritier, H. Mattiat, A.-K. Thamm, R. Schirhagl, et al., Physical Review Applied 2021, 15, 2 L021001.
- [7] C. Reinhardt, T. Müller, A. Bourassa, J. C. Sankey, Physical Review X 2016, 6, 2 021001.
@@
- [8] S. S. Verbridge, J. M. Parpia, R. B. Reichenbach, L. M. Bellan, H. G. Craighead, Journal of Applied Physics 2006, 99, 12 124304.
- [9] R. A. Norte, J. P. Moura, S. Gröblacher, Physical review letters 2016, 116, 14 147202.
@@
- [10] A. H. Ghadimi, S. A. Fedorov, N. J. Engelsen, M. J. Bereyhi, R. Schilling, D. J. Wilson, T. J. Kippenberg, Science 2018, 360, 6390 764.
- [11] S. A. Fedorov, A. Beccari, N. J. Engelsen, T. J. Kippenberg, Physical Review Letters 2020, 124, 2 025502.
@@
- [12] M. J. Bereyhi, A. Arabmoheghi, A. Beccari, S. A. Fedorov, G. Huang, T. J. Kippenberg, N. J. Engelsen, Physical Review X 2022, 12, 2 021036.
@@
- [13] A. Beccari, D. A. Visani, S. A. Fedorov, M. J. Bereyhi, V. Boureau, N. J. Engelsen, T. J. Kippenberg, Nature Physics 2022, 18, 4 436.
@@
- [14] E. Romero, V. M. Valenzuela, A. R. Kermany, L. Sementilli, F. Iacopi, W. P. Bowen, Physical Review Applied 2020, 13, 4 044007.
- [15] Y. S. Klaß, Ph.D. thesis, Technische Universität München, 2022.
@@
- [16] S. K. Manjeshwar, A. Ciers, F. Hellman, J. Bläsing, A. Strittmater, W. Wieczorek, arXiv preprint arXiv:2211.12469 2022.
- [17] X. W. Gu, Z. Wu, Y.-W. Zhang, D. J. Srolovitz, J. R. Greer, Nano letters 2013, 13, 11 5703.
- [18] X. Zhao, B. Mao, M. Liu, J. Cao, S. J. Haigh, D. G. Papageorgiou, Z. Li, R. J. Young, Advanced Functional Materials 2022, 32, 42 2202373.
@@
- [19] M. J. Bereyhi, A. Beccari, S. A. Fedorov, A. H. Ghadimi, R. Schilling, D. J. Wilson, N. J. Engelsen, T. J. Kippenberg, Nano letters 2019, 19, 4 2329.
- [20] G. Kwon, H.-H. Jo, S. Lim, C. Shin, H.-H. Jin, J. Kwon, G.-M. Sun, Journal of materials science 2015, 50 8104.
- [21] H. I. Rasool, C. Ophus, W. S. Klug, A. Zettl, J. K. Gimzewski, Nature communications 2013, 4, 1 2811.
- [22] T. Zhang, X. Li, H. Gao, International Journal of Fracture 2015, 196 1.
- [23] M. Goldsche, J. Sonntag, T. Khodkov, G. J. Verbiest, S. Reichardt, C. Neumann, T. Ouaj, N. von den Driesch, D. Buca, C. Stampfer, Nano letters 2018, 18, 3 1707.
- [24] C. Lee, X. Wei, J. W. Kysar, J. Hone, science 2008, 321, 5887 385.
- [25] M. Wang, C. Yan, L. Ma, N. Hu, M. Chen, Computational Materials Science 2012, 54 236.
- [26] J. Xu, G. Yuan, Q. Zhu, J. Wang, S. Tang, L. Gao, ACS nano 2018, 12, 5 4529.
- [27] R. Qu, M. Calin, J. Eckert, Z. Zhang, Scripta Materialia 2012, 66, 10 733.
- [28] H. Jiang, T. Shang, H. Xian, B. Sun, Q. Zhang, Q. Yu, H. Bai, L. Gu, W. Wang, Small Structures 2021, 2, 2 2000057.
- [29] A. L. Greer, Science 1995, 267, 5206 1947.
- [30] M. Telford, Materials today 2004, 7, 3 36.
@@
- [31] M. Wijesundara, R. Azevedo, Silicon carbide microsystems for harsh environments, volume 22, Springer Science & Business Media, 2011.
- [32] R. Gerhardt, Properties and applications of silicon carbide, BoD–Books on Demand, 2011.
@@
- [33] T. Kimoto, J. A. Cooper, Fundamentals of Silicon Carbide Technology: Growth, Characterization, Devices and Applications, Wiley-IEEE Press, 2014.
- [34] B. Morana, G. Pandraud, J. Creemer, P. Sarro, Materials Chemistry and Physics 2013, 139, 2-3 654.
@@
- [35] C. Iliescu, D. P. Poenar, In Physics and Technology of Silicon Carbide Devices. IntechOpen, 2012.
- [36] T. Blum, B. Dresler, M. Hoffmann, et al., Surface and Coatings Technology 1999, 116 1024.
- [37] L. Jiang, X. Chen, X. Wang, L. Xu, F. Stubhan, K.-H. Merkel, Thin Solid Films 1999, 352, 1-2 97.
- [38] S. Castelletto, A. Boretti, Journal of Physics: Photonics 2020, 2, 2 022001.
@@
- [39] M. Köhler, M. Pomaska, P. Procel, R. Santbergen, A. Zamchiy, B. Macco, A. Lambertz, W. Duan, P. Cao, B. Klingebiel, et al., Nature Energy 2021, 6, 5 529.
- [40] L. Sementilli, E. Romero, W. P. Bowen, Advanced Functional Materials 2022, 32, 3 2105247.
@@
- [41] T.-K. Nguyen, H.-P. Phan, H. Kamble, R. Vadivelu, T. Dinh, A. Iacopi, G. Walker, L. Hold, N.-T. Nguyen, D. V. Dao, ACS applied materials & interfaces 2017, 9, 48 41641.
@@
- [42] H. A. Atwater, A. R. Davoyan, O. Ilic, D. Jariwala, M. C. Sherrott, C. M. Went, W. S. Whitney, J. Wong, Nature materials 2018, 17, 10 861.
- [43] B. Morana, Ph.D. thesis, TU Delft, 2015.
- [44] J. Chu, D. Zhang, Journal of Micromechanics and Microengineering 2009, 19, 9 095020.
- [45] M. Imran, M. Mahendran, P. Keerthan, Journal of Constructional Steel Research 2018, 143 131.
- [46] Y. Zhou, Y. Wang, P. Mallick, Materials Science and Engineering: A 2004, 381, 1-2 355.
- [47] K.-S. Chen, A. Ayon, S. M. Spearing, Journal of the American Ceramic Society 2000, 83, 6 1476.
- [48] J. Cui, Z. Zhang, H. Jiang, D. Liu, L. Zou, X. Guo, Y. Lu, I. P. Parkin, D. Guo, ACS nano 2019, 13, 7 7483.
- [49] D. J. Shuman, A. L. Costa, M. S. Andrade, Materials characterization 2007, 58, 4 380.
- [50] J.-H. Kim, S.-C. Yeon, Y.-K. Jeon, J.-G. Kim, Y.-H. Kim, Sensors and Actuators A: Physical 2003, 108, 1-3 20.
- [51] Y. S. Klaß, J. Doster, M. Bückle, R. Braive, E. M. Weig, Applied Physics Letters 2022, 121, 8 083501.